OpenStudy (anonymous):

Given a circle with a radius of 6, what is the length of an arc measuring 60°? A.1/2pi B. 2pi c. 3/2pi D. 3pi

OpenStudy (dumbcow):

convert 60 degrees to radians 60 * pi/180 = pi/3 length = angle*radius = (pi/3)*6 = 2pi